# Time-bin entanglement in the deterministic generation of linear photonic cluster states

### v2

@@ III Stabilizer Generator Expectation Values
In Fig. 6, we calculate stabilizer generator expressions for up to the 15th expression. Notably, these stabilizers maintain a consistent behavior across a spectrum of loss conditions. This consistency is crucial, as it demonstrates the potential for creating large cluster states; diminishing effects from losses do not appear to accumulate through the protocol cycle. This stability plays a critical role in the practicality of developing complex cluster and graph states, even within lossy environments. Additionally, it proves advantageous for the numerical analysis that only the first three stabilizer generator expectation values need to be assessed, since the subsequent values resemble the initial ones almost exactly.
@@
$\mathcal{W}=(N-1)-\langle X^{(1)}Z^{(2)}\rangle-\sum_{i>1}\langle Z^{(i-1)}X^{(i)}Z^{(i+1)}\rangle\,.$ (12)
@@
Here, we use pulse configurations that maximize the stabilizer generator expectation values for $\Phi=X$. Evaluating Eq. 12, we determine conditions for entanglement by asserting that $\mathcal{W}<0$. Assuming uniformity among all stabilizer generator expectation values - with the exception of the initial $\langle XZ\rangle$ we deduce an expression for the lower bound of the length $N$ of the linear cluster state. We omit the superscript indices from the expressions under the assumption that the expectation values are identical for all indices, which results in
$N<\frac{\langle XZ\rangle}{1-\langle ZXZ\rangle}+1\,.$ (13)
@@
Applying Eq. 13 to the three cases presented in Fig. 6, we round down to the nearest integer, resulting in the following length estimates for linear cluster states that can be generated with the investigated scheme: $N_{\text{no loss}}<98$, indicating the potential size without any loss mechanisms; $N_{\text{Radiative}}<9$, reflecting the impact of radiative losses; and $N_{\text{Dephasing}}<3$, showcasing the limitations imposed by spin dephasing of the ground states, which needs to be minimized for the generation of larger cluster states. We display the maximum achievable length of the linear cluster state for a given stabilizer magnitude according to Eq. 13 in Fig. 6. These results again support that slight enhancements (reductions) in rotation fidelity lead to disproportionately large extensions (decline) in the achievable lengths of cluster states, a phenomenon only perceptible through the lens of stabilizer generator expectation values. While minor percentage improvements in rotation fidelity might not visibly enhance the protocol, our analysis of the stabilizers reveals their profound impact.
@@
We find the indistinguishability of the single photons generated in the system to be inherently high, a direct consequence of the photon emission originating from a modified two-level system [72, 73]. This resemblance of a two-level system ensures that each emitted photon possesses uniform phase and spectral characteristics, critical factors that enhance their indistinguishability [74]. In quantum information processing, the ability to produce highly indistinguishable photons is paramount, as it facilitates efficient quantum interference, a cornerstone for quantum computing algorithms and photonic quantum information protocols. When dealing with indistinguishable photons, it is also possible to expand the foundational cluster state into multidimensional states through specific measurements [18].
@@ IV Conclusion
We conducted an in-depth study of the generation of time-bin entangled photons using a $\Lambda$-type electronic system, relevant for various physical realizations of emitters. Our microscopic numerical investigation reveals that conventional metrics used to analyze excitation-emission protocols, such as fidelities of quantum-state rotations [68], fall short of fully assessing the usefulness of the correlated quantum states generated, especially when faced with imperfections and losses in a realistic setting. Even when high rotation fidelities are achieved, quantum correlations generated can be surprisingly low, diminishing the usefulness of the generated quantum state. In the present study we use expectation values of stabilizer generators as a more robust measure of correlations in a multi-photon quantum state. These include full sensitivity to dephasing and losses, providing a more nuanced understanding of the system’s quantum dynamics. For the generation of linear photonic cluster states, we find that even minor rates of spin dephasing (that hardly affect protocol fidelities) can result in a very significant reduction of stabilizer generator magnitudes and, with it, reduction of quantum correlations.
Combining these insights we explicitly numerically demonstrate the generation of large linear photonic cluster states comprised of a two-digit number of photonic qubits, illustrating the need for precise control over system losses as well as timing and phases of laser pulses used for excitation and gating. In addition to high degrees of quantum correlations achieved, emitted photons also show high degrees of indistinguishability, essential for subsequent quantum interference measurements. Further increase of quantum correlations for given loss parameters could potentially be achieved by reducing lengths of time bins for example by use of optical cavities for the cycling transitions, improving scalability of the generation of correlated multi-photon states for quantum information processing purposes.
@@ Acknowledgements.
This work was supported by the Deutsche Forschungsgemeinschaft (German Research Foundation) through the transregional collaborative research center TRR142/3-2022 (231447078, project C09), the Photonic Quantum Computing initiative (PhoQC) of the state ministry (Ministerium fur Kultur und Wissenschaft des Landes Nordrhein-Westfalen), and with computing time provided by the Paderborn Center for Parallel Computing, PC2. We greatfully acknowledge fruitful discussions with Stefan Schulz and Gediminas Juska from Tyndall National Institute.
Appendix A Correlation Functions
@@
A.1 General Second and Third Order Correlation Functions
@@
using $\left\{EE,EL,LE,LL\right\}$ as the basis. Analogously, we define the three photon density matrix using $\mathcal{G}^{(3)}(t_{1},t_{2},t_{3})$ in Eq. 4, which corresponds to the basis using all of the available three-way permutations of $E$ and $L$. While these would, in theory, suffice for the evaluation of two and three photon entanglement, scalability becomes an obvious problem for larger photon numbers.
A.2 Examples of Correlation Functions

### v3

@@ III Stabilizer Generator Expectation Values
In Fig. 6, we calculate stabilizer generators for up to the 15th expression. Notably, these stabilizers maintain a consistent behavior across a spectrum of loss conditions, meaning all stabilizers remain close to their mean value. In other words, the standard deviation of all of the stabilizers of a given linear cluster state is small. This consistency is crucial, as it demonstrates the potential for creating large cluster states; diminishing effects from losses do not appear to accumulate through the protocol cycle. The stabilizers remain constant because the loss between the first and second $Z$ projections within each stabilizer expression is consistent, making the starting time of the stabilizer evaluation insignificant. This stability plays a critical role in the practicality of developing complex cluster and graph states, even within lossy environments. Additionally, it proves advantageous for the numerical analysis that only the first three stabilizer generator expectation values need to be assessed, since due to their local nature the subsequent values resemble the initial ones almost exactly.
@@
$\langle\mathcal{W}\rangle=(N-1)-\langle\hat{X}^{(1)}\hat{Z}^{(2)}\rangle-\sum_{i>1}\langle\hat{Z}^{(i-1)}\hat{X}^{(i)}\hat{Z}^{(i+1)}\rangle\,.$ (12)
@@
Here, we use pulse configurations that maximize the stabilizer generator expectation values for $\Phi=X$. Evaluating Eq. 12, we determine conditions for entanglement by asserting that $\langle\mathcal{W}\rangle<0$. Assuming uniformity among all stabilizer generator expectation values - with the exception of the initial $\langle XZ\rangle$ we deduce an expression for the lower bound of the length $N$ of the linear cluster state. We omit the superscript indices from the expressions under the assumption that the expectation values are identical for all indices, which results in
$N<\frac{\langle\hat{X}\hat{Z}\rangle}{1-\langle\hat{Z}\hat{X}\hat{Z}\rangle}+1\,.$ (13)
@@
Applying Eq. 13 to the three cases presented in Fig. 6, we round down to the nearest integer, resulting in the following length estimates for linear cluster states that can be generated with the investigated scheme: $N_{\text{no loss}}<98$, indicating the potential size without any loss mechanisms (here only limited by the finite pulse lengths and corresponding bandwiths); $N_{\text{Radiative}}<9$, reflecting the impact of radiative losses; and $N_{\text{Dephasing}}<3$, showcasing the limitations imposed by spin dephasing of the ground states, which needs to be minimized for the generation of larger cluster states. We display the minimum achievable length of the linear cluster state for a given stabilizer magnitude according to Eq. 13 in Fig. 6. These results again support that slight enhancements (reductions) in rotation fidelity lead to disproportionately large extensions (decline) in the achievable lengths of cluster states, a phenomenon only perceptible in a detailed analysis of quantum correlations in the generated multi-partite photon state, here performed through the lens of stabilizer generator expectation values. While minor percentage improvements in rotation fidelity might not visibly enhance the protocol, our analysis of the stabilizers reveals their profound impact.
@@
Although not visualized in this work, our simulations reveal the indistinguishability of the single photons generated in the system to be inherently high, a direct consequence of the photon emission originating from a modified two-level system [73, 74]. Consequently, the indistinguishability is lowered for non-zero spin dephasing of the ground states, again closely mirroring an isolated two level system. This resemblance of a two-level system ensures that each emitted photon possesses uniform phase and spectral characteristics, critical factors that enhance their indistinguishability [75]. In quantum information processing, the ability to produce highly indistinguishable photons is paramount, as it facilitates efficient quantum interference, a cornerstone for quantum computing algorithms and photonic quantum information protocols. When dealing with indistinguishable photons, it is also possible to expand the foundational cluster state into multidimensional states through specific measurements [18]. For the case with strong radiative losses and dephasing in Fig. 6, the indistinguishability and single-photon purity of the qubit states each stay above $95\%$, approaching unity values when losses approach zero.
@@ IV Conclusion
We conducted an in-depth theoretical study of the generation of time-bin entangled photons using a $\Lambda$-type electronic system, relevant for various physical realizations of emitters. Specifically, we focused our research on a system present in quantum dot molecule composed of two individual quantum dot emitters, resulting in a deterministic source. Our microscopic numerical investigation reveals that conventional metrics used to analyze excitation-emission protocols, such as fidelities of quantum-state rotations [68], fall short of fully assessing the usefulness of the correlated quantum states generated, especially when faced with imperfections and losses in a realistic setting. Even when high rotation fidelities are achieved, quantum correlations generated can be surprisingly low, diminishing the usefulness of the generated quantum state. In the present study we use expectation values of stabilizer generators as a more robust measure of correlations in a multi-photon quantum state. These include full sensitivity to dephasing and losses, providing a more nuanced understanding of the system’s quantum dynamics and usefulness as a quantum resource. For the generation of linear photonic cluster states, we find that even minor rates of spin dephasing (that hardly affect protocol fidelities) can result in a very significant reduction of stabilizer generator magnitudes and, with it, reduction of quantum correlations.
Combining these insights we explicitly numerically demonstrate the generation of large linear photonic cluster states comprised of a two-digit number of photonic qubits. These results require optimized timing and phases of laser pulses used for excitation and gating. Notably, maximizing the rotation fidelities towards unity values, where precision is crucial, appears as a requirement for the generation of large numbers of entangled qubits. In addition to high degrees of quantum correlations achieved, photons emitted in our calculations also show inherently high degrees of indistinguishability due to the emission resulting from an effective two level system. High photon indistinguishability is a property essential for subsequent quantum interference measurements. Further increase of quantum correlations for given loss parameters could potentially be achieved by reducing lengths of time bins for example by use of optical cavities for the cycling transitions, improving scalability of the generation of correlated multi-photon states for quantum information processing purposes. Furthermore, by calculating even higher order correlation functions, the evaluation of different, more complicated cluster states can be done, albeit with significantly increased numerical effort.
@@ Acknowledgements.
This work was supported by the Deutsche Forschungsgemeinschaft (German Research Foundation) through the transregional collaborative research center TRR142/3-2022 (231447078, project C09), the Photonic Quantum Computing initiative (PhoQC) of the state ministry (Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen), and with computing time provided by the Paderborn Center for Parallel Computing, PC2. We greatfully acknowledge fruitful discussions with Stefan Schulz and Gediminas Juska from Tyndall National Institute and with Zahra Raissi from Paderborn University.
Appendix A Hamiltonian & Equation of Motion
We use the following Hamiltonian to describe the system numerically:
$H$ $=H_{0}+H_{I}$ (14)
$H_{0}$ $=\sum_{\begin{subarray}{c}o\in\{\text{G}_{1},\text{G}_{2},\text{X}_{1},\text{X}_{2},\text{T},\text{U}\}\end{subarray}}E_{o}|o\rangle\langle o|$ (15)
[…]
With generally non-zero ground state energies $E_{\text{G}_{i}}$, excited state energies $E_{\text{X}_{i}}$ as well as target and unwanted state energies $E_{\text{T}}$ and $E_{\text{U}}$. Here, $\Omega_{\text{cyc}}(t)=\Omega_{i,\text{X}_{i}}(t)$, exciting any of the two ground states $\text{G}_{i}$ into their respective excited states $\text{X}_{i}$. The target and unwanted transitions are given by $\Omega_{1,2}(t)=\Omega_{i,\text{U}}(t)+\Omega_{i,\text{T}}(t)$, with $i\in\{1,2\}$. The numerical evaluation is performed using a variable time step Runge Kutta solver, solving the von Neumann equation
[…]
in the interaction frame for $H_{0}$ with Lindbad rates
$\mathcal{L}^{\hat{O}}[\rho(t)]=\sum_{\hat{O}}\frac{1}{2}\left(2\hat{O}\rho\hat{O}^{\dagger}-\hat{O}^{\dagger}\hat{O}\rho-\rho\hat{O}^{\dagger}\hat{O}\right)$ (18)
for the radiative loss ($\hat{O}=\sqrt{\gamma^{\text{R}}}|i\rangle\langle j|$) and spin dephasing ($\hat{O}=\sqrt{\gamma^{\text{D}}}\left(|j\rangle\langle j|-|i\rangle\langle i|\right)$ where $E_{i}<E_{j}$). Compare with [35, 36] for further details.
Appendix B Correlation Functions
@@
B.1 General Second and Third Order Correlation Functions
@@
using $\left\{EE,EL,LE,LL\right\}$ as the basis. In this case, the index ordering corresponds to the operator ordering. Analogously, we define the three photon density matrix using $\mathcal{G}^{(3)}(t_{1},t_{2},t_{3})$ in Eq. 4, which corresponds to the basis using all of the available three-way permutations of $E$ and $L$. While these would, in theory, suffice for the evaluation of two and three photon entanglement, scalability becomes an obvious problem for larger photon numbers.
B.2 Examples of Correlation Functions